# X-ray free electron laser SPI single-particle diffractive imaging

### v1

@@ 5.2 Blind case
The quantitative results are listed in Table 2. The SSP methods again achieve the best image metric results of all compared methods and recovers over 90% of the positions correctly, but regarding position recovery shows no advantage over using only a H-TV prior. This suggests that the score-based prior is helpful for retrieving better images, but not necessarily for easing the overall position-blind reconstruction problem. RED-Diff and our SSP variant with position deltas perform significantly worse at recovering the positions correctly at only 52-62% correctly recovered. Since RED-Diff also treats the fitted positions as fixed values rather than distributions (see Section 3), we conclude that the added randomness from fitting a random distribution on the positions is very helpful for position recovery. Nonetheless, both methods reach fair image metric values, suggesting that using only a subset of the measurements may be enough to retrieve a decent image in this specific simulated measurement setup.
@@ 5.2.1 Importance of the probe structure
Without a phase mask and $d_{\text{ap}}\in\{\nicefrac{{1}}{{2}},\nicefrac{{1}}{{4}}\}$, we recover almost no positions correctly and fail to produce usable images. While $d_{\text{ap}}=\nicefrac{{1}}{{8}}$ recovers 89% of positions correctly, its final images (Fig. 4(a), center) are contaminated with low-frequency artifacts and the image metrics are subpar, which may be explained by the lack of high-frequency structure in the probe due to the low-pass from the small aperture. In comparison, adding block-wise random phase masks to the previously unusable aperture with $d_{\text{ap}}=\nicefrac{{1}}{{2}}$ leads to the best image quality we observed and also improves the position recovery, with a 25.3 dB aPSNR and 95% of positions recovered correctly for the block size $b=4$. We further note that $b=8$ still performs rather well compared to $b=4$ despite illuminating a significantly smaller region of the object (see Fig. 2), again suggesting that probe structure is more helpful than probe size for our task.
@@ 5.2.2 Different levels of measurement noise
Here we analyze and compare the behavior of the blind SSP method, blind RED-Diff, and blind VI without an image prior when increasing the level (scale) of the Gaussian measurement noise $\sigma_{\varepsilon}$ from the default $\sigma_{\varepsilon}=0.005$. We show quantitative results in Table 4, where we also list the corresponding measurement SNR, and qualitative results for the SSP method in Fig. 6(a), where we also show the fitted per-pixel variance of the variational Gaussian that can inform uncertainty estimation. We can observe that (1) the data-driven priors allow for at least some image and position recovery even under -9.5 dB measurement noise, where the VI method without an image prior completely fails; (2) the SSP method performs best and most reliably under all noise levels and, surprisingly, recovers 69% of positions correctly even at the highest noise level; (3) the per-pixel uncertainty from SSP (Fig. 6) is, at least at the highest noise level, informative about regions that are heavily affected by artifacts. For a complete visual comparison of all methods under all measurement noise levels, see Fig. 13. We note here that the $\lambda_{\text{RD}}$ parameter of RED-Diff could be tuned further in dependence of $\sigma_{\varepsilon}$ in order to potentially improve the results, but we do not follow this here.
@@ 5.2.4 Phase-only objects
We therefore added a log-barrier prior loss on the positions as described in Section 4.2, with the weight empirically set to $100$. With this, we found a reliable reconstruction ability, with all objects and repeated runs leading to satisfactory image quality. In Fig. 8, we show successful example reconstructions for the standard probe $d_{\text{ap}}=\nicefrac{{1}}{{2}},b=4$ with the blind SSP method. Nonetheless, in terms of metrics, we find a worsened cRMS value of 0.06 ± 0.06 and a posCorrect value of only 61.60 ± 14.04 (cf. Table 2), showing the increased difficulty of this problem, particularly for position recovery.
@@ 5.3 Towards uncertainty quantification
Figure 10: The per-pixel uncertainty without a prior and with a surrogate score-based prior (SSP), either estimated from the fitted diagonal covariance of the Gaussian distribution $q_{\chi}$ (Fitted) or estimated from a population of three independently fitted means (Population). For comparison, we also show the per-pixel to the ground-truth object, i.e., |xx^||x-$\hat${x}| (Error).
@@
6 Summary and Outlook
In this work, we investigate the novel blind inverse problem of position-blind ptychography for the first time, with possible applications in biological single-particle diffractive imaging (SPI) and ptychographic imaging under extreme movement of the sample. To attack this problem, we develop and demonstrate a Bayesian variational inference approach that can employ modern data-driven image priors and classic model-driven priors, and compare the developed method against another method from recent literature for data-driven solutions to blind inverse problems.
We evaluate our approach for increasingly difficult scenarios including phase-only objects, measurements with Poisson noise, and weak-phase phase-only objects under the presence of a beamstop, and show that we can reliably achieve successful reconstructions in all but the most difficult scenario. We investigate the underlying reasons for the difficulty of variants of our position-blind ptychography imaging problem, and propose remedies in the form of structured illumination and additional prior terms on the positions.
@@ 6 Summary and Outlook
Since we treat only a simplified 2-D variant of the full 3-D position- and rotation-blind ptychography problem that would arise in SPI, we explicitly list our simplifying assumptions for clarity. We expect that it is possible for future works to build upon our Bayesian approach to include scenarios close to real-world SPI experiments towards 3D images with unknown positions and unknown rotations, unstable illumination (unknown or varying probe), and structured measurement noise, which we hope can eventually unlock this new imaging modality for real-world biological structure investigations.
@@ Acknowledgments
MB, LK and TR acknowledge support from DESY (Hamburg, Germany), a member of the Helmholtz Association HGF. This research was supported in part through the Maxwell computational resources operated at Deutsches Elektronen-Synchrotron DESY, Hamburg, Germany. This work was carried out while LK was a member of DESY. MB, TG, LK, TR and SW acknowledge funding by the German Federal Ministry of Research, Technology and Space (BMFTR) under grant agreement No. 01IS24072A (COMFORT). MB, LK and TR acknowledge partial funding by the DAAD project 57698811 ”Bayesian Computations for Large-scale (Nonlinear) Inverse Problems in Imaging”.
Appendix A RED-diff Algorithm for Blind Inverse Problems

### v2

@@ 5.2 Blind case
The quantitative results are listed in Table 3. The SSP methods again achieve the best image metric results of all compared methods and recovers more than 94% of the positions correctly on average, but regarding position recovery shows no advantage over using only a H-TV prior. This suggests that the score-based prior is helpful for retrieving better images, but not necessarily for easing the overall position-blind reconstruction problem. RED-Diff and our SSP variant with position deltas perform significantly worse at recovering the positions correctly at only 53-65% correctly recovered. Since RED-Diff also treats the fitted positions as fixed values rather than distributions (see Section 3), we conclude that the added randomness from fitting a random distribution on the positions is very helpful for position recovery. Nonetheless, both methods reach fair image metric values, suggesting that using only a subset of the measurements may be enough to retrieve a decent image in this specific simulated measurement setup.
@@ 5.2.1 Importance of the probe structure
Without a phase mask and $d_{\text{ap}}\in\{\nicefrac{{1}}{{2}},\nicefrac{{1}}{{4}}\}$, we recover almost no positions correctly and fail to produce usable images. While $d_{\text{ap}}=\nicefrac{{1}}{{8}}$ recovers 89% of positions correctly, its final images (Fig. 4(a), center) are contaminated with low-frequency artifacts and the image metrics are subpar, which may be explained by the lack of high-frequency structure in the probe due to the low-pass from the small aperture. In comparison, adding block-wise random phase masks to the previously unusable aperture with $d_{\text{ap}}=\nicefrac{{1}}{{2}}$ leads to the best image quality we observed and also improves the position recovery, with a 25.3 dB aPSNR and 94% of positions recovered correctly for the block size $b=4$, see Table 4. We further note that $b=8$ still performs rather well compared to $b=4$ despite illuminating a significantly smaller region of the object (see Fig. 2), again suggesting that probe structure is more helpful than probe size for our task.
@@
5.2.2 Convergence behavior and runtime efficiency
We now turn our attention to the numerical convergence behavior of the compared methods in the blind and non-blind settings. We treat the negative log-likelihood (NLL), $\log p(y|x)$, as the quantity that should ideally converge. We show convergence curves, i.e., the NLL plotted against the iteration number, for each compared method on a single example image and reconstruction run in Fig. 6. We compare against the NLL of the ground-truth image and ground-truth positions relative to the simulated noisy measurements $\{y_{k}\}$, which we call ground-truth NLL.
Non-blind setting
In the non-blind setting, see Fig. 6 on the left, it can be seen that most methods converge below the ground-truth NLL, indicating overfitting to the measurement noise. The only methods that avoid this behavior and converge with values close to the ground-truth NLL are VI (SSP), VI with the Huber-TV prior, and RED-Diff, indicating that the priors and choice of prior weighting used in these three methods successfully combat the overfitting. Interestingly, RED-Diff only reaches values close to the ground-truth NLL close to the end of the reconstruction. We suspect that this is due to the fixed reverse diffusion schedule that RED-Diff follows [77], which makes the generative prior focus only on the low image frequencies at first and only uses the full image frequency content during the final iterations. In contrast, SSP uses all diffusion timesteps through weighted random sampling throughout the reconstruction, see (3.15). For all methods besides RED-Diff, the convergence plots suggest that the full 10,000 iterations are not needed in this non-blind setting, and one could terminate at much fewer iterations.
Blind setting
In the blind setting, see Fig. 6 on the right, none of the methods converge to a value close to the ground-truth NLL. This is expected, as the blind problem is of higher difficulty and not all positions are recovered correctly by any method (cf. Table 3). In contrast to the non-blind setting, the plotted curves demonstrate that all methods profit from the full 10,000 iterations here, and may even improve the NLL further if given more iterations.
The SSP method reaches the lowest values, followed by VI with a Huber-TV prior, VI with no prior, and the non-variational optimization method with no prior. For the two methods without a prior, the unfavorable aPSNR and cRMS values (cf. Table 3) suggest that the low NLL values are due to overfitting to measurement noise. This is also reflected in the reconstructed images, see Fig. 3(b). Furthermore, the plot for the prior-free optimization-based method during early iterations suggests a more unstable and complex behavior than for the other methods. RED-Diff and the optimization method with a Huber-TV prior exhibit a rather large gap in NLL to the other methods. We attribute this also to the comparably worse position recovery observed in Fig. 3(b).
Figure 6: Evolution of the negative log-likelihood along the optimization trajectory for different methods. The dashed line indicates $\log p(y|x^{*})$, where $x^{*}$ denotes the ground truth image.
Related to the convergence behavior, the runtime and memory efficiency of the compared methods is also of practical interest. In Table 5, we show and compare the empirical runtime and peak memory usage of each used method variant. It can be seen that the VI-based methods all take more than 4 hours for the reconstruction, while the optimization-based methods and RED-Diff take around one hour. The memory usage of the VI-based methods is also substantially higher. The reason for this lies mostly in the use of a batch size $B>1$ to perform parallel evaluation of the likelihood and prior terms on multiple image and position samples from the variational distributions, see Section 3.1.4. We support this idea by an additional line in Table 5 when using $B=1$ instead. Evidently, the GPU does not effectively parallelize over the batch, since the runtime increases by a factor very close to $B=4$. The peak GPU memory is increased by a factor of around 3.2 compared to $B=1$. Interestingly, the use of a diffusion model as in VI (SSP) and RED-Diff does not significantly increase the runtime nor memory usage, which suggests that both may be dominated by the costly repeated evaluation of the likelihood on $N=100$ high-dimensional diffraction pattern measurements. Note that a runtime- or memory-optimal implementation is out of scope for this work, and we expect that improvements to the runtime and GPU memory usage can be made in future work.
Method Runtime GPU Mem.
VI (SSP) 4:49 h 19082 MiB
VI (SSP) with $B=1$ 1:11 h 5986 MiB
VI (H-TV prior) 4:15 h 17124 MiB
VI (No prior) 4:15 h 17122 MiB
RED-Diff 1:11 h 6712 MiB
Opt. (H-TV prior) 1:04 h 6479 MiB
Opt. (No prior) 0:55 h 6472 MiB
Table 5: Overall runtime in hours and peak GPU memory usage for one image reconstruction, using each method. Determined on an NVIDIA A100 80GB GPU.
@@ 5.2.3 Different levels of measurement noise
Here we analyze and compare the behavior of the blind SSP method, blind RED-Diff, and blind VI without an image prior when increasing the level (scale) of the Gaussian measurement noise $\sigma_{\varepsilon}$ from the default $\sigma_{\varepsilon}=0.005$. We show quantitative results in Table 6, where we also list the corresponding measurement SNR, and qualitative results for the SSP method in Fig. 7(a), where we also show the fitted per-pixel variance of the variational Gaussian that can inform uncertainty estimation. We can observe that (1) the data-driven priors allow for at least some image and position recovery even under -9.5 dB measurement noise, where the VI method without an image prior completely fails; (2) the SSP method performs best and most reliably under all noise levels and, surprisingly, recovers 69% of positions correctly on average even at the highest noise level; (3) the per-pixel uncertainty from SSP (Fig. 7) is, at least at the highest noise level, informative about regions that are heavily affected by artifacts. A complete visual comparison of the methods under all measurement noise levels is shown in Fig. 16. We note here that the $\lambda_{\text{RD}}$ parameter of RED-Diff could be tuned further in dependence of $\sigma_{\varepsilon}$ in order to potentially improve the results, but we do not follow this here.
@@ 5.2.5 Phase-only objects
We therefore added a log-barrier prior loss on the positions as described in Section 4.2, with the weight empirically set to $100$. With this, we found a reliable reconstruction ability, with all objects and repeated runs leading to satisfactory image quality. In Fig. 9, we show successful example reconstructions for the standard probe $d_{\text{ap}}=\nicefrac{{1}}{{2}},b=4$ with the blind SSP method. Nonetheless, in terms of metrics, here we find a worsened cRMS value of 0.06 ± 0.06 and a posCorrect value of only 61.60 ± 14.04, which are both worse than the values for blind reconstructions of complex objects that are also absorbing, see Table 3. This shows the increased difficulty of this problem variant, particularly for position recovery.
@@ 5.3 Towards uncertainty quantification
Figure 11: The per-pixel uncertainty without a prior and with a surrogate score-based prior (SSP), either estimated from the fitted diagonal covariance of the Gaussian distribution $q_{\chi}$ (Fitted) or estimated from a population of three independently fitted means (Population). Note that all reconstructions are obtained by VI. For comparison, we also show the per-pixel to the ground-truth object, i.e., |xx^||x-$\hat${x}| (Error).
@@
6 Discussion
In the following we discuss possible paths towards reconstruction in more realistic measurement setups. Furthermore, we highlight the risks and challenges connected to reconstruction methods based on generative models.
6.1 Towards a more realistic SPI setting
@@
In this paper we treat only a simplified 2-D variant of the full 3-D position- and rotation-blind ptychography problem that would arise in SPI. We expect that it is possible for future works to build upon our Bayesian approach to include scenarios closer to real-world SPI experiments, towards 3D images with unknown positions and unknown rotations, unstable illumination (unknown or varying probe), and structured measurement noise. In the following we comment on how we hope to eventually unlock this new imaging modality for real-world biological structure investigations.
6.1.1 On Assumption 2.1
Figure 12: Reconstructions under a mismatched probe function, using variational inference with a score-based prior (top row) or no prior (bottom row). $f_{\text{Zern}}$ denotes the factor applied to the Zernike polynomial coefficients for the probe used during reconstruction, compared to the original value of 5 used for the simulated measurements.
Assuming that the probe is known, as in Assumption 2.1, is not generally sensible in real ptychographic measurements [87]. A relaxation of this assumption within our framework should be possible by treating the entire probe function as another parameter to recover, given that the scheme remains reasonably stable under an imperfectly estimated probe.
As a motivating example, we briefly investigate the effect of a mismatched probe function on reconstructions by our VI scheme. First, we use the same simulated configuration to generate the same measurements as in Section 5.2. Then, we simulate three mismatched probes by increasing the magnitude of the Zernike aberration coefficients from the “true” factor of 5 that was used to generate the measurements, to factors 6, 7, and 8. We keep the random phase mask the same, assuming that it can be characterized well and is stable. We then run one reconstruction each for each of these mismatched probes. We show the results in Fig. 12. While the scheme recovers a worse and more blurry image, as well as lower PSNR and fewer correct positions – as one would expect from using the wrong probe function – it still succeeds to recover a reasonable image in the overall position-blind ptychography problem. Furthermore, interestingly, the use of the generative score-based surrogate prior (SSP) results in more reliable position recovery in this probe-mismatched setting than using no prior. Empirically, therefore, we argue that the scheme Eq. 3.19 is stable enough under a probe-mismatched estimated measurement operator to allow adding the probe optimization as a third alternating (or joint) optimization step in Eq. 3.19. We leave detailed investigations of this to future work.
6.1.2 On Assumption 2.3
Imaging three-dimensional objects with ptychographic methods has been successfully achieved in various works, and we refer the reader to [87, Sec. 17.6] for a detailed overview. Rodenburg & Maiden [87] note two major types of approaches: Ptycho-tomography and multislice ptychography. Multislice ptychography assumes that the sample is not rotated between measurements, and is hence not applicable to the general SPI scenario, see Section 6.1.3.
Ptycho-tomography has been successfully realized both computationally and experimentally in several prior works [31, 53, 14, 6, 87]. However, these works generally assume that scan positions and object rotations are known, which likely presents a substantially easier task than the fully position- and rotation-blind one. While ptycho-tomography typically cannot account for multiple scattering [87], this should not present a problem for X-ray SPI of biomolecules, due to the very small overall sample thickness of even large proteins, where the first-order Born approximation (cf. [16]) should hold.
Performing ptycho-tomography with generative priors has also previously been successfully realized [6]. In this work, the authors also briefly suggest that future work can add a position correction to their ADMM-based scheme, by introducing additional auxiliary variables. Similarly, we expect that by using a forward model similar to [6], introducing learnable rotational parameters to our VI scheme as described in the next Section 6.1.3, treating the imaged target object $x$ as a 3-dimensional array, and training appropriate three-dimensional generative priors, a future extension to three-dimensional objects is possible. Using three-dimensional diffusion-based priors for multi-slice ptychographic reconstructions has already been successfully performed in [68]. Nonetheless, the difficulty of the fully blind problem variant, and any possibility of convergence, is unknown and remains a highly interesting future direction.
6.1.3 On Assumption 2.4
In Assumption 2.4, we assumed that the object is always oriented the same in every measurement, and no rotational parameters must be recovered. This assumption generally does not hold in real SPI settings, where the macromolecules under investigation enter the interaction region with random orientations (and translations, which we however already treat). To some extent, this problem can potentially be addressed experimentally by physically aligning macromolecules such as proteins, e.g., with lasers [3]. However, research on this is still ongoing, and as [3] notes, the unknown molecule orientation is so far treated in silico instead. Our variational approach can be extended in this direction, by also modeling and optimizing the distribution of all per-image object rotation angles $\{\varphi_{k}\}_{k=1,\ldots,K}$ (or in 3D, e.g., quaternions), though care must be taken to use distributions that can correctly model the involved circular statistics. The distribution parameters would then also be optimized through backpropagation with respect to the loss (3.17). This is viable with existing software, e.g., the kornia Python package offers a differentiable rotate function. Whether it is better to treat this optimization as part of a single update step for the positions and rotations, or to perform further splitting of the alternating minimization (3.19), is a topic reserved for future work.
6.1.4 On Assumption 2.5
In Assumption 2.5, we assume that all diffraction patterns used for reconstruction correspond to probe positions where the probe illuminated the object significantly (no measurements of empty space). This assumption is in line with prior literature on SPI, see for instance [96] where patterns were classified as hits or non-hits by a simple thresholding of the total measurement intensity. Such a classification can also be straightforwardly used as a data preprocessing step for our methods.
6.2 Risk of generative models for scientific applications
Figure 13: Position-blind ptychography reconstructions from our variational inference methods, of an image from a human face dataset which is mismatched to the training dataset of the score-based prior, see Section 4.3. The columns indicate different choices of prior, the rows indicate increasing levels of measurement noise. In the most noisy setting, the learned surrogate score-based prior (SSP) hallucinates building-like structures (mouth) and overly smooths other regions such as the eyes – both features that are likely under the learned prior, rather than reflect the actual ground-truth image.
Generative models are promising new tools in the field of inverse problems and, as also shown in the results presented herein, can aid with successful reconstructions, especially under strong measurement noise. However, as discussed in various prior works on data-driven reconstruction methods [17, 5, 80, 49, 66], they also carry risks with regards to the trustworthiness expected of scientific practice. For instance, in image regions that are not well-described by the measurement data, generative priors tend to dominate the reconstructed image, which can result in image features that are not meaningfully supported by the measurements and may thus not correspond to physical reality – a phenomenon which, in an anthropomorphizing manner, is often referred to as hallucination [49, 80]. Such hallucinations are clearly a problem for scientific inquiry and the reliability of results, as they can lead to false and unreliable interpretations. As an extreme example, when used for protein structure investigations, a generative model could produce an entire protein substructure that the real protein does not have, leading to entirely false conclusions and catastrophic downstream decisions, e.g. for drug development.
We note, however, that even simple model-based priors such as TV, in principle suffer from errors introduced by said priors, which is inherent to the Bayesian formulation. In an image region that is not captured well by the measurements, the TV prior can dominate, which would result in a uniformly flat region in the reconstruction even if this does not correspond to the imaged object. Nonetheless, the much higher simplicity and interpretability of such model-based priors suggests a higher level of trust since it is considered to be easier to spot and analyze when and why such faulty reconstructions arise.
The problem of so-called hallucinations becomes particularly clear when the generative prior is trained on image data that has distinctly different features than the images captured of the desired real physical objects, which is often referred to as data mismatch. To empirically demonstrate that this problem does arise in our context, we run a reconstruction on an image from a human face dataset with different statistics (CelebA-HQ [62] at 256x256 resolution, converted to grayscale and used directly for both amplitude and phase without any smoothing) while using the score model trained on our simulated and smoothed INRIA data, see Section 4.3. We show the results in Fig. 13. We can observe the different types of artifacts introduced by different prior choices. In the no prior case the reconstruction tends to overfit to the noise. As expected the TV creates flat regions in the image. For the SSP prior, we do not observe similar artifacts. Instead the generative model places features into the reconstruction that are comparable to the ones from the INRIA dataset.
6.3 Summary and Outlook
In this work, we investigate the novel blind inverse problem of position-blind ptychography for the first time, with possible applications in biological single-particle diffractive imaging (SPI) and ptychographic imaging under extreme movement of the sample. To attack this problem, we derive and demonstrate a Bayesian variational inference approach that can employ both modern data-driven image priors and classic model-driven priors, and compare this method against non-variational optimization-based approaches and another method from recent literature for data-driven solutions to blind inverse problems, RED-Diff [77, 1].
We evaluate our approach for increasingly difficult scenarios including phase-only objects, measurements with Poisson noise, and weak-phase phase-only objects under the presence of a beamstop, and show that we can reliably achieve successful reconstructions in all but the most difficult scenario. We investigate the underlying reasons for the difficulty of variants of our position-blind ptychography imaging problem, and propose remedies in the form of structured illumination and additional prior terms on the positions. We highlight all assumptions underlying our simplified position-blind ptychography problem, and offer guidance for future extensions towards realistic single-particle diffractive imaging (SPI) settings. Overall, this work contributes a new interesting inverse problem to the literature and provides a solid foundation for future investigations into this new imaging problem. We hope that the ideas and observations presented here can spark future investigations in this field.
@@ Acknowledgments
MB, LK and TR acknowledge support from DESY (Hamburg, Germany), a member of the Helmholtz Association HGF. This research was supported in part through the Maxwell computational resources operated at Deutsches Elektronen-Synchrotron DESY, Hamburg, Germany. This work was carried out while LK and TR were members of DESY. MB, TG, LK, TR and SW acknowledge funding by the German Federal Ministry of Research, Technology and Space (BMFTR) under grant agreement No. 01IS24072A (COMFORT). MB, LK and TR acknowledge partial funding by the DAAD project 57698811 ”Bayesian Computations for Large-scale (Nonlinear) Inverse Problems in Imaging”.
[Uncaptioned image][Uncaptioned image]
Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon Europe research and innovation programme (grant agreement No. 101198055, project acronym NEITALG).
Appendix A Additional figures
Figure 14: Comparison of VI reconstruction with the Huber-TV prior under different prior weights $\lambda\in{5,10,20}$.
Figure 15: All probes and corresponding apertures used for evaluation in this work, in particular when comparing the effect of different probes. Our default choice throughout the work is the probe with $d_{\text{ap}}=\nicefrac{{1}}{{2}},b=4$, which is visually duplicated here for easier visual comparison and a cleaner layout. $b=0$ indicates no phase mask was used for the respective probe.
Figure 16: Reconstructed images of a test object from three methods (SSP, RED-Diff, VI without an image prior) under three increasing levels of Gaussian measurement noise $\sigma_{\varepsilon}$.
Appendix B RED-diff Algorithm for Blind Inverse Problems